### v1

# Bounding the Treewidth of Outer $k$-Planar Graphs via Triangulations

Abstract
The treewidth is a structural parameter that measures the tree-likeness of a graph. Many algorithmic and combinatorial results are expressed in terms of the treewidth. In this paper, we study the treewidth of outer $k$-planar graphs, that is, graphs that admit a straight-line drawing where all the vertices lie on a circle, and every edge is crossed by at most $k$ other edges.
Wood and Telle [New York J. Math., 2007] showed that every outer $k$-planar graph has treewidth at most $113k+$ 11 using so-called planar decompositions, and later, Auer et al. [Algorithmica, 2016] proved that the treewidth of outer $1$-planar graphs is at most $3$, which is tight.
In this paper, we improve the general upper bound to $21.5k+$ 2 and give a tight bound of $4$ for $k=2$. We also establish a lower bound: we show that, for every even $k$, there is an outer $k$-planar graph with treewidth $k+2$. Our new bound immediately implies a better bound on the cop number, which answers an open question of Durocher et al. [GD 2023] in the affirmative.
Our treewidth bound relies on a new and simple triangulation method for outer $k$-planar graphs that yields few crossings with graph edges per edge of the triangulation. Our method also enables us to obtain a tight upper bound of $k+2$ for the separation number of outer $k$-planar graphs, improving an upper bound of $32k+$ 3 by Chaplick et al. [GD 2017]. We also consider outer min-$k$-planar graphs, a generalization of outer $k$-planar graphs, where we achieve smaller improvements.
keywords:
treewidth, outer $k$-planar graphs, outer min-$k$-planar graphs, cop number, separation number
1 Introduction
Treewidth measures the tree-likeness of a graph via so-called tree decompositions. A tree decomposition of a graph $G$ covers the vertex set of $G$ by bags such that every edge is in some bag and the bags form a tree such that, for every vertex $v$ of $G$, the bags that contain $v$ form a subtree. The width of a tree decomposition is the maximum size of a bag minus one. The treewidth of $G$ is the minimum width over all tree decompositions of $G$.
Treewidth is an important structural parameter because the running time of many graph algorithms and algorithms for drawing graphs depend on the treewidth. Recently treewidth-based techniques have been successfully applied for convex drawings, that is, straight-line drawings where vertices lie on a circle. For example, Bannister and Eppstein [8] showed that if a graph admits a convex drawing with at most $k$ crossings in total, then its treewidth is bounded by a function of $k$. Furthermore, they used this fact to give, for a fixed $k$, a linear-time algorithm (via extended monadic second-order logic and Courcelle’s theorem [13]) to decide whether a convex drawing with at most $k$ crossings exists for a given graph. Chaplick et al. [12] generalized this to bundled crossings (a crossing of two bundles of edges such that in each bundle the edges travel in parallel, which is counted as one crossing). Another prominent class of graphs with convex drawings is the class of outer $k$-planar graphs, that is, graphs that admit a convex drawing where each edge is crossed at most $k$ times. Unlike $k$-planar graphs (without the restriction on the placement of vertices), the treewidth of outer $k$-planar graphs can be bounded by a function of $k$ only (see discussion below). Similarly to Bannister and Eppstein [8], Chaplick et al. [10] used this to test in linear time, for any fixed $k$, whether a given graph is full outer $k$-planar, i.e., whether it admits an outer $k$-planar drawing where no crossing appears on the boundary of the outer face.
For disambiguation, recall that $k$-outerplanar graphs are defined as follows. A drawing of a graph is $1$-outerplanar if all vertices of the graph lie on the outer face. For $k>1$, a drawing is $k$-outerplanar if deleting the vertices on the outer face yields a $(k-1)$-outerplanar drawing. For $k\geq 1$, a graph is $k$-outerplanar if it admits a $k$-outerplanar drawing.
In this paper, we are particularly interested in the treewidth of outer $k$-planar graphs. Note that every graph is outer $k$-planar for some value of $k$. For a graph $G$, let $\operatorname{lcr}^{\circ}(G)$ denote the convex local crossing number of $G$ – the smallest $k$ such that $G$ is outer $k$-planar. Consult Schaefer’s survey [19] for details on this and many other types of crossing numbers. Outer $k$-planar graphs admit balanced separators of size $O(k)$ (more below) associated with the drawing, which makes it possible to test, for any fixed $k$, outer $k$-planarity in quasi-polynomial time [10]. This implies that the recognition problem is not NP-hard unless the Exponential Time Hypothesis fails [10].
We also consider a generalization of outer $k$-planar graphs, namely outer min-$k$-planar graphs, which are graphs that admit a convex drawing where, for every pair of crossing edges, at least one edge is crossed at most $k$ times. Wood and Telle [20, Proposition 8.5] showed that any (min-) outer $k$-planar graph has treewidth at most $113k+$ 11. More precisely, they showed that, for every outer min-$k$-planar graph $G$, one obtains a tree decomposition of width at most $113k+$ 11 from a planar decomposition of $G$, which is a generalization of a tree decomposition. The proof uses a result by Bodlaender [9] on $k$-outerplanar graphs.
For constant $k$, better treewidth bounds are known. A folklore result is that outerplanar graphs ($k=0$) have treewidth at most $2$. Auer et al. [5, 6] showed that maximal outer $1$-planar graphs are chordal graphs and, therefore, have treewidth at most $3$, which is tight.
Another parameter linked to the treewidth is the separation number of a graph. A pair of vertex sets $(A,B)$ is a separation of a graph $G$ if $A\cup B=V(G)$ and there is no edge between $A\setminus B$ and $B\setminus A$. A separation $(A,B)$ is said to be balanced if the sizes of both $A\setminus B$ and $B\setminus A$ are at most $32n/$ 3, where $n$ is $|V(G)|$. For a (balanced) separation $(A,B)$, the set $A\cap B$ is called (balanced) separator, and $|A\cap B|$ is the order of $(A,B)$. The separation number of $G$, $\operatorname{sn}(G)$, is the minimum integer $k$ such that every subgraph of $G$ has a balanced separation of order $k$. Note that, for any subgraph $G^{\prime}$ of $G$, $\operatorname{tw}(G^{\prime})\leq\operatorname{tw}(G)$. This implies that, for every graph $G$, $\operatorname{sn}(G)\leq\operatorname{tw}(G)+1$ [15]. On the other hand, Dvořák and Norin [15] showed that, for every graph $G$, $\operatorname{tw}(G)\leq 15\operatorname{sn}(G)$.
Our contribution.
Given an outer $k$-planar drawing $\Gamma$ of a graph $G$, let the outer cycle of $G$ be the cycle that connects the vertices of $G$ in the order along the circle on which they lie in $\Gamma$ (even if $G$ does not contain all edges of this cycle). We introduce two simple methods to construct, given an outer $k$-planar drawing $\Gamma$ of a graph $G$, a triangulation of the outer cycle of $G$ with the property that each edge of the triangulation is crossed by at most $k$ edges of $G$; see Section 3. The resulting triangulations yield the following bounds; see Section 4.
  • We improve the upper bound of Wood and Telle [20] regarding the treewidth of outer $k$-planar graphs from $113k+$ 11 to $21.5k+$ 2 (Theorem 4.4 in Section 4.1). Our proof is constructive and implies a practical algorithm; the tree decomposition that we obtain follows the weak dual of the triangulation. For outer $2$-planar graphs, our methods yield an upper bound of 4 (Theorem 4.3), which is tight due to $K_{5}$.
  • Chaplick et al. [10] showed that, for every outer $k$-planar graph $G$, its separation number is at most $32k+$ 3. We improve this upper bound to $k+2$; see Section 4.2.
  • We give new lower bounds of $k+2$ for both the treewidth and the separation number of outer $k$-planar graphs; see Section 5. Note that the latter bound is tight (Theorem 5.1).
  • Durocher et al. [14] recently proved that the cop number of general 1-planar graphs is not bounded, but the maximal 1-planar graphs have cop number at most 3. Since it is known [16] that for every graph $G$, $\operatorname{cop}(G)\leq\operatorname{tw}(G)/2+1$, Durocher et al. [14] observed that the treewidth bound of Wood and Telle yields, for every outer $k$-planar graph $G$, that $\operatorname{cop}(G)\leq 1.5k+6.5$. They asked explicitly whether the multiplicative factor of $1.5$ can be improved. We answer this question in the affirmative since our new treewidth bound immediately yields the better bound $\operatorname{cop}(G)\leq 0.75k+1.75$.
  • For outer min-$k$-planar graphs, our triangulation improves the bound of Wood and Telle $113k+$ 11 slightly to $13k+$ 1 and gives a bound $12k+$ 1 on the separation number; see Section 4.
Related results.
A structurally similar type of result is known for $2$-layer $k$-planar graphs. A $2$-layer $k$-planar graph is a bipartite graph that admits a $k$-planar straight-line drawing with vertices placed on two parallel lines. The maximal $2$-layer 00-planar graphs (called caterpillars) are the maximal pathwidth-1 graphs [17, 18]. Angelini et al. [2] showed that $2$-layer $k$-planar graphs have pathwidth at most $k+1$ and gave a lower bound of $(k+3)/2$ for every odd number $k$. Hence, their result can be considered as a smooth generalization of the caterpillar result. Similarly, our treewidth bound $21.5k+$ 2 for outer $k$-planar graphs smoothly generalizes the treewidth-$2$ bound for maximal outerplanar graphs.
2 Preliminaries
A convex drawing $\Gamma$ of a graph is a straight-line drawing where the vertices of the graph are placed on different points of a circle, which we call the circle of $\Gamma$. An outer $k$-planar drawing $\Gamma$ of a graph is a convex drawing such that every edge crosses at most $k$ other edges. A counterclockwise walk of the circle of $\Gamma$ yields a cyclic order on the vertices of the graph. We say that two sets on distinct four vertices $\{v_{1},w_{1}\}$, and $\{v_{2},w_{2}\}$ are intertwined in $\Gamma$ if they are ordered $\langle v_{1},v_{2},w_{1},w_{2}\rangle$ or $\langle v_{1},w_{2},v_{2},w_{1}\rangle$ in this cyclic order. Observe that two edges $\{v_{1},w_{1}\}$ and $\{v_{2},w_{2}\}$ cross in $\Gamma$ if and only if they span four different vertices and are intertwined. In the remainder of this section, we make some simple observations that will be helpful later; they can be skipped by more experienced readers. {observation} Let $G$ be a graph. If $\Gamma$ is an outer $k$-planar drawing of $G$, then every convex drawing $\Gamma^{\prime}$ of $G$ with the same cyclic vertex order as in $\Gamma$ is also an outer $k$-planar drawing of $G$.
Proof.
As the cyclic orders on vertices of $G$ defined by $\Gamma$ and $\Gamma^{\prime}$ are the same, exactly the same pairs of edges cross in $\Gamma$ and in $\Gamma^{\prime}$. ∎
In the definition of convex drawings, we could have allowed placing vertices on arbitrary convex shapes (instead of a circle) and drawing edges as curves (instead of straight-line segments) – as long as every curve is drawn inside the circle and no two curves cross more than once. Using curves will be sometimes convenient in our proofs later.
{observation}
Let $G$ be a graph, let $\Gamma$ be an outer $k$-planar drawing of $G$, and let $c$ be a curve that joins vertices $v$ and $w$ of $G$ and is contained inside the circle of $\Gamma$. If $c$ crosses at most $l$ edges of $\Gamma$, then the straight-line segment $\overline{vw}$ also crosses at most $l$ edges of $\Gamma$.
Proof.
As $c$ must cross every edge $\{v^{\prime},w^{\prime}\}$ of $G$ that is intertwined with $\{v,w\}$, we get that $c$ must have at least one crossing with every edge crossed by the straight-line segment $\overline{vw}$. ∎
A graph $G$ is maximal outer $k$-planar if $G$ is outer $k$-planar and $G$ does not contain any vertex pair $e\in V^{2}(G)\setminus E(G)$ such that the graph $(V(G),E(G)\cup\{e\})$ is still outer $k$-planar. Since removing edges increases neither treewidth nor separation number, we are interested in properties of maximal outer $k$-planar graphs.
{observation}
If $G$ is a maximal outer $k$-planar graph with at least three vertices, then, in every outer $k$-planar drawing of $G$, the outer face is bounded by a simple cycle.
Proof.
Let $v$ and $w$ be two vertices that are consecutive in the cyclic order defined by some outer $k$-planar drawing $\Gamma$ of $G$. For a contradiction, suppose that $v$ and $w$ are not adjacent. As $v$ and $w$ are consecutive on the circle of $\Gamma$, the set $\{v,w\}$ is not intertwined with any edge in $\Gamma$. Thus, $\Gamma$ can be extended to an outer $k$-planar drawing of $G+\{v,w\}$, which contradicts the maximality of $G$. ∎
We remark that Sections 2, 2 and 2 analogously hold for outer min-$k$-planar graphs.
3 Triangulations
In this section, we present two simple strategies to triangulate maximal outer $k$-planar graphs. These triangulations will serve as tools to construct tree decompositions and balanced separators in Section 4. We assume that some outer $k$-planar drawing $\Gamma$ of a graph $G$ on $n$ vertices is given. During the triangulation procedures, we will modify the drawing, but we will never change the cyclic order of the vertices. Thus, let $v_{1},v_{2},\dots,v_{n}$ be the vertices of $G$ in the cyclic (counterclockwise) order defined by $\Gamma$. As we focus on maximal graphs, by Section 2, we have that $\langle v_{1},v_{2},\dots,v_{n},v_{1}\rangle$ is a cycle in $G$. Our goal is to construct a triangulation $\mathcal{T}$ of this cycle such that the edges of $\mathcal{T}$ cross only a limited number of edges of $G$. We refer to the edges of $\mathcal{T}$ as links in order to distinguish them from the edges of $G$. Every edge of the outer cycle is a link in $\mathcal{T}$; we call these links outer links. We will select a set of $n-3$ other pairwise non-intertwined pairs of vertices as inner links that, together with the outer links, form a triangulated $n$-gon. Note that some of the inner links may coincide with edges of $G$. We say that a link is pierced by an edge of $G$ if their endpoints are intertwined in the cyclic order. The piercing number of a link is the number of edges of $G$ that pierce the link. In particular the piercing number of outer links is 0, and if a link coincides with an edge, then its piercing number is at most $k$ by the outer $k$-planarity of $\Gamma$. We define the edge piercing number of $\mathcal{T}$ to be the maximum piercing number of any link of $\mathcal{T}$. In Section 4 we show how to use triangulations with small edge piercing number.
Our splitting procedure starts by declaring the link $\{v_{1},v_{n}\}$ active and considers all other vertices to lie on the right side of the link. See Fig. 1 for illustration. Clearly, the link $\{v_{1},v_{n}\}$ is not pierced. In each recursive step, the input is an active link $\{v_{i},v_{k}\}$ with $1\leq i<k\leq n$ (promised to be pierced at most some limited number of times), along with a distinguished right side $v_{i+1},\dots,v_{k-1}$, where currently no inner links have been selected apart from $\{v_{i},v_{k}\}$. (Vertices $v_{1},\dots,v_{i-1},v_{k+1},\dots,v_{n}$ form the left side.) The goal is to pick, among the vertices on the right side, a split vertex $v_{j}$ such that the new links $\{v_{i},v_{j}\}$ and $\{v_{j},v_{k}\}$ are pierced by at most some limited number of edges. The new links are added to $\mathcal{T}$ (completing the triangle $v_{i}v_{j}v_{k}$) and become active. The link $\{v_{i},v_{k}\}$ ceases to be active. The split gives rise to two new splitting instances $\{v_{i},v_{j}\}$ and $\{v_{j},v_{k}\}$, which are then solved recursively.
(a) input graph
(b) step 1
(c) step 2
(d) step 3
(e) triangulation
Figure 1: Steps of the splitting procedure. The active edge is purple, the (new/old) links are (dark/light) blue, the left/right side of the active edge is red/green. The split vertex is big.
The set of edges of $G$ that pierce the active link are the piercing edges. Note that every piercing edge has one left and one right endpoint. Intuitively, the edges with two left endpoints are of no concern to the splitting procedure. The next lemma allows us to push all crossings among piercing edges to the left of $\{v_{i},v_{k}\}$.
Lemma 3.1.
Given an outer $k$-planar drawing $\Gamma$ of a graph $G$ and an active link $\{v_{i},v_{k}\}$, there exists an outer $k$-planar drawing $\Gamma^{\prime}$ with the same cyclic order as $\Gamma$, and all crossings among the edges piercing the active link are drawn to the left of the active link.
Proof 3.2.
We draw the vertices of $G$ on the unit circle, keeping the cyclic order defined by $\Gamma$. We set $\varepsilon>0$ to a sufficiently small value. Then we place $v_{i}$ at $(0,-1)$, $v_{k}$ at $(0,1)$, all the vertices from the left side of the link within arc distance $\varepsilon$ of $(-1,0)$ and spread the vertices from the right side of the link evenly along the right semicircle, see Fig. 2.
Figure 2: We can push crossings among the edges that pierce the active link $\{v_{i},v_{k}\}$ to the left. We show only the top- and bottommost endpoints on the left. Crossings occur exclusively in the green area, which can be made arbitrarily small.
Now, Section 2 yields that the new drawing is an outer $k$-planar drawing of $G$ with the same cyclic order. As every pair of intertwined piercing edges has their left endpoints in arc distance at most $\varepsilon$ from $(-1,0)$ and their right endpoints in arc distance at least $\pi/n$ from each other, the crossing of two such segments is to the left of the active link.
Lemma 3.1 gives us an equivalent drawing of $G$ in which piercing edges cross the active link in the same order from bottom to top as their right endpoints in the cyclic order. This allows us to draw the piercing edges without crossings on the right side of the active link.
Now, we are ready to present our first and basic triangulation method. Asymptotically, it yields a worse bound than Lemma 3.5, but it serves as a gentle introduction into our techniques, and it will be used for small values of $k$ and for outer min-$k$-planar graphs later.
Lemma 3.3.
For $k\geq 1$, every outer $k$-planar drawing of a maximal outer $k$-planar graph admits a triangulation of the outer cycle with edge piercing number at most $12k-$ 1.
Proof 3.4.
We prove the lemma by providing an appropriate splitting procedure that recursively constructs the triangulation. Let $\Gamma$ be an outer $k$-planar drawing of $G$. We prove by induction on the number of steps of the splitting procedure that every active link has piercing number at most $12k-$ 1. The claim is true in the beginning, as we start with the link $\{v_{1},v_{n}\}$, which is not pierced at all.
Now we consider a later step of the procedure. Let $\lambda=\{v_{i},v_{k}\}$ denote the active link. For brevity, set $u=v_{i}$ and $v=v_{k}$. There are two cases.
In the first case, $\lambda$ is not pierced by any edge of $G$. Observe that $\{v_{i},v_{i+1}\}$ is an edge, so $u$ has at least one neighbor in the right side of $\lambda$ in $G$. We select the split vertex, $w$, to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$ in the circular order; see Fig. 3(a). In this case, the link $\{u,w\}$ coincides with an edge of $G$, and thus is pierced by at most $k$ edges. To see that the link $\{w,v\}$ is also pierced by at most $k$ edges, let $c^{\prime\prime}$ be a curve that starts in $w$, goes along $\{w,u\}$ towards $u$ on the right side (in the direction of walking), then follows $\{u,v\}$ on the right side, and ends at $v$. This curve crosses only the edges that pierce the link $\{u,w\}$, so at most $k$ edges pierce the link $\{w,v\}$ by Section 2.
In the second case, $\lambda$ is pierced by edges of $G$. Using Lemma 3.1, we modify the drawing $\Gamma$ so that the piercing edges cross $\lambda$ in the same bottom-to-top order as their right endpoints. Let $e_{1},e_{2},\dots,e_{\ell}$ denote the piercing edges and $w_{1},w_{2},\dots,w_{\ell}$ be their respective, not necessarily different, right endpoints in the order from bottom to top. By induction, we have $\ell\leq 2k-1$. Then we pick the middle edge $e_{j}$, where $j=\lceil(\ell+1)/2\rceil$, and set the split vertex $w=w_{j}$.
(a) constructing the new triangle $\triangle$ if $\lambda$ is not pierced by any other edge of $G$.
(b) constructing the new triangle $\triangle$ bounded by $\lambda$, $c^{\prime}$, and $c^{\prime\prime}$; the orange edges that pierce both $c^{\prime}$ and $c^{\prime\prime}$ also cross the edge $e_{j}$.
Figure 3: Constructing a triangulation $\mathcal{T}$ with edge piercing number $12k-$ 1.
To bound the number of edges that pierce the link $\{u,w\}$, we draw the curve $c^{\prime}$ as in Fig. 3(b). We start the curve $c^{\prime}$ at $u$ and follow $\lambda$ until before it crosses $e_{j}$, then follow $e_{j}$ until we arrive at $w$. The curve $c^{\prime}$ has crossings with at most $\lceil(\ell-1)/2\rceil$ edges that pierce $\lambda$, and at most $k$ edges that have a crossing with $e_{j}$. Thus, there are at most $\lceil(\ell-1)/2\rceil+k\leq\lceil(2k-2)/2\rceil+k=2k-1$ edges that pierce the link $\{u,w\}$. We can argue symmetrically for the link $\{w,v\}$; see curve $c^{\prime\prime}$ in Fig. 3(b).
As we have bounded the piercing number of every active link during the triangulation by $12k-$ 1, we get the desired bound of $12k-$ 1 for the edge piercing number of the constructed triangulation.
Next, we present our refined triangulation method, where we consider an additional edge to get better bounds for the edge piercing number.
Lemma 3.5.
For every $k\geq 1$, every outer $k$-planar drawing of a maximal outer $k$-planar graph admits a triangulation of the outer cycle with edge piercing number at most $k$.
Proof 3.6.
The proof is similar to the proof of Lemma 3.3 and we use the same terminology. We present a splitting procedure and show that every active link is pierced by at most $k$ edges. For the base of induction, the link $\{v_{1},v_{n}\}$ has no piercing edges. Now, let $\lambda=\{v_{i},v_{k}\}$ be the active link and set $u=v_{i}$ and $v=v_{k}$. If $\lambda$ has no piercing edges, we proceed as in the proof of Lemma 3.3 and pick the split vertex to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$. Otherwise, $\lambda$ is pierced by edges of $G$. Using Lemma 3.1, we assume that the piercing edges cross the active link in the same bottom-to-top order as their right endpoints. Now, let $e_{1},e_{2},\ldots,e_{\ell}$ be the piercing edges. We pick the middle edge $e_{j}$ with $j=\lceil(\ell+1)/2\rceil$. If $e_{j}$ has no more crossings on the right of the active link, then we pick the right endpoint of $e_{j}$ to be the split vertex. Both curves that start in the right endpoint of $e_{j}$, follow $e_{j}$ to $e$, and then follow $\lambda$ to either $u$ or $v$ have at most $\lceil(\ell-1)/2\rceil\leq k$ crossings each, so we can proceed. Now, let $x$ be the first crossing on the edge $e_{j}$ that occurs on the right side of $\lambda$, and let $\hat{e}$ denote the edge that crosses $e_{j}$ at $x$; see Fig. 4(a).
We first consider the case where $\hat{e}$ and $\lambda$ are disjoint. Let $a$ and $b$ be the numbers of crossings on $\hat{e}$ on the two sides of $x$. As $a+b+1\leq k$, we get that $\min\{a,b\}\leq\lfloor(k-1)/2\rfloor$. We select the split vertex $w$ to be the endpoint of $\hat{e}$ on the side that has fewer, i.e., $\min\{a,b\}$, crossings along $\hat{e}$. We now assume $w_{j}<w<v$. The case $u<w<w_{j}$ is symmetric.
(a) $\lambda$ and $\hat{e}$ do not share an endpoint; the new triangle $\triangle$ is bounded by $\lambda$, $c^{\prime}$, and $c^{\prime\prime}$.
(b) $\lambda$ and $\hat{e}$ share an endpoint; the new triangle $\triangle$ is bounded by $\lambda$, $\hat{e}$, and $c^{\prime\prime}$.
Figure 4: Constructing a triangulation $\mathcal{T}$ with edge piercing number $k$.
We bound the piercing number of the link $\{v,w\}$. We start the curve $c^{\prime\prime}$ at $v$ and follow the link $\lambda$ until just before it crosses $e_{j}$, then follow $e_{j}$ until just before it crosses $\hat{e}$, then follow $\hat{e}$ until we arrive at $w$; see Fig. 4(a). In the first part, the curve crosses at most $\lceil(\ell-1)/2\rceil\leq\lceil(k-1)/2\rceil$ edges. The curve does not cross any edge in the middle part, as the crossing between $e_{j}$ and $\hat{e}$ is the first crossing along $e_{j}$ starting from $\lambda$. In the last part, the curve crosses at most $\min\{a,b\}\leq\lfloor(k-1)/2\rfloor$ edges. Thus, by Section 2, the piercing number of the link $\{v,w\}$ is at most $\lceil(k-1)/2\rceil+\lfloor(k-1)/2\rfloor\leq k-1$.
To bound the piercing number of the link $\{u,w\}$, we start the curve $c^{\prime}$ at $u$ and follow the link $\lambda$ until just after it crosses $e_{j}$ and then follow the curve $c^{\prime\prime}$ to $w$, see Fig. 4(a). This time, the number of crossings on the curve $c^{\prime}$ is at most $k$. Thus, the piercing number of the link $\{u,w\}$ is at most $k$.
Now we consider the last case, where $\hat{e}$ and $\lambda$ have a common vertex. In this case, we can choose the other vertex of $\hat{e}$ as $w$; see Fig. 4(b). Without loss of generality, we assume that $u$ is the common vertex. The edge $\{u,w\}$ has at most $k$ crossings by $k$-planarity. To bound the piercing number of the link $\{v,w\}$, observe that the edge $\{u,w\}$ has exactly $j$ crossings on the curve from $v$ to $x$ (including $x$). Therefore, as we did in the above, the edge $\{v,w\}$ is crossed by at most $(\ell-j)+(k-j)=k-(2j-\ell)$ edges. Note that $2j-\ell$ is $1$ if $\ell$ is $odd$ and 00 otherwise.
The piercing number of a triangle of $\mathcal{T}$ is the sum of the piercing numbers of the links that form the sides of the triangle. We define the triangle piercing number of $\mathcal{T}$ to be the maximum piercing number of any triangle of $\mathcal{T}$. Note that the triangle piercing number of a triangulation is at most three times its edge piercing number, but it can be smaller. If $k$ is odd, we obtain a slightly stronger result; If $k$ is odd, for every triangle $t$ that we create in the proof of Lemma 3.5, at least one edge of $t$ has at most $k-1$ crossings.
Lemma 3.7.
For every odd $k\geq 1$, every outer $k$-planar drawing of a maximal outer $k$-planar graph admits a triangulation of the outer cycle with edge piercing number at most $k$ and triangle piercing number at most $13k-$ 1.
Proof 3.8.
Following the proof of Lemma 3.5, in all cases, there is an edge with piercing number at most $k-1$. If $\lambda$ has no piercing edges, its piercing number is $10\leq k-$ 1. Otherwise, if $\hat{e}$ and $\{u,v\}$ are disjoint, $\{v,w\}$ is pierced by at most $k-1$ edges. In the last case, where $\hat{e}$ and $\{u,v\}$ have a common endpoint, $\{v,w\}$ is pierced by at most $k-1$ edges if $\ell$ is odd. If $\ell$ is even, it implies $\ell\neq k$ and therefore the link $\lambda$ is pierced by $\ell\leq k-1$ edges.
Slightly improving the analysis in the proof of Lemma 3.5 for $k=2$ yields the following.
Lemma 3.9.
Every outer $2$-planar drawing of a maximal outer $2$-planar graph admits a triangulation of the outer cycle with edge piercing number at most $2$ and triangle piercing number at most $4$.
Proof 3.10.
We bound the triangle piercing number more carefully. If $\lambda$ has no piercing edges or $e_{j}$ has no crossings on the right side, then the claim is clearly true. If $e_{j}$ has a crossing, the curves $c^{\prime}$ and $c^{\prime\prime}$ meet each other near $e_{j}$ with exactly $l$ crossings in total and then proceed to $w$ without crossings. Hence, the triangle piercing number is at most $4$.
Lastly, we combine the two strategies from Lemmas 3.3 and 3.5 to obtain a triangulation method for outer min-$k$-planar graphs.
Lemma 3.11.
For every $k\geq 1$, every outer min-$k$-planar drawing of a maximal outer min-$k$-planar graph admits a triangulation of the outer cycle with edge piercing number at most $12k-$ 1 and triangle piercing number at most $36k-$ 3.
Proof 3.12.
We proceed as in the proof of Lemma 3.3. We want to give a splitting procedure for which every active link has piercing number at most $12k-$ 1. Consult Fig. 3 for a reminder of the notation and the strategy. As $\Gamma$ is now an outer min-$k$-planar drawing, edges may have more than $k$ crossings. We call such edges heavy. Other edges, i.e., edges with no more than $k$ crossings, are light. By definition, heavy edges cross only light edges.
The first thing to notice is that as we allow for heavy edges now, the previous proof breaks in both cases when the curves $c^{\prime}$, $c^{\prime\prime}$ are routed along heavy edges. In particular, in the first case, $\lambda$ has no piercing edges, and we select the split vertex $w$ to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$. If edge $\{u,w\}$ is heavy, then we cannot bound the number of crossings along the curve $c^{\prime\prime}$. In the second case, we select the split vertex to be the right endpoint $w_{j}$ of the middle piercing edge $e_{j}$. If edge $e_{j}$ is heavy, then we cannot bound the numbers of crossings along $c^{\prime}$ and $c^{\prime\prime}$.
To resolve the issue with the second case, we apply the strategy from the proof of Lemma 3.5. Observe that if $e_{j}$ is heavy, then for the first crossing $x$ with an edge $\hat{e}$ on the right side of $\lambda$, $\hat{e}$ is a light edge. Similarly, as in the proof of Lemma 3.5, see Fig. 4, we select the split vertex $w$ to be the endpoint of $\hat{e}$ that is different than $u$ and $v$ and in case both endpoints are, the one on the side of $x$ with fewer crossings. The piercing numbers of new links are at most $\lceil(2k-1)/2\rceil+k-1\leq 2k-1$.
To resolve the issue with the first case, we use a similar strategy. When edge $\{u,w\}$ is heavy, let $x$ to be the first crossing on $\{u,w\}$ from $u$ with an edge $\hat{e}$. Notice that $\hat{e}$ is light. We select the split vertex $w$ to be any right endpoint of $\hat{e}$. The curves that start at $w$, go along $\hat{e}$ to $x$, then along $\{u,w\}$ to $u$ and optionally to $v$ along $\lambda$ have at most $k$ crossings, only with edges that also cross $\hat{e}$.
Remark 3.13.
Each step of the splitting procedures used in Lemmas 3.33.11 can be implemented to run in $O(k)$ time, and the corresponding triangulations can be constructed in $O(nk)$ time.
4 Applications
In this section, we show several applications of our triangulations from Section 3. Among others, we improve the bound of Wood and Telle [20] on the treewidth of outer $k$-planar graphs from $113k+$ 11 to $21.5k+$ 2 (see Section 4.1) and improve the bound of Chaplick et al. [10] on the separation number of outer $k$-planar graphs from $32k+$ 3 to $k+2$; see Section 4.2. In addition, we improve the bound of Wood and Telle of $113k+$ 11 on the treewidth of outer min-$k$-planar graphs to $13k+$ 1; see Theorem 4.6.
4.1 Treewidth
Now we show our main tool to obtain a better upper bound on the treewidth. Note that we do not refer specifically to outer $k$-planar graphs since our tool can be used for any drawing of a graph whose vertices lie on the outer face.
Consider a graph $G$ with a drawing $\Gamma$ whose outer face is bounded by a simple cycle $C$ that contains all vertices of $G$. Let $\mathcal{T}$ be a planar triangulation of $C$ and $f$ be some triangular face of $\mathcal{T}$. We call an edge of $E(G)\setminus E(\mathcal{T})$ short with respect to $f$ if one of its endpoints is on $f$ and it pierces a link of $f$. We call an edge of $E(G)\setminus E(\mathcal{T})$ long with respect to $f$ if neither of its endpoints is on $f$ and it pierces two links of $f$.
Lemma 4.1.
If $G$ is a graph with a convex drawing whose outer cycle admits a triangulation $\mathcal{T}$ with triangle piercing number at most $c$, then $\operatorname{tw}(G)\leq(c+5)/2$.
Proof 4.2.
Let $T$ be a tree obtained by taking the weak dual graph of $\mathcal{T}$ and associating each node of $T$ with a bag that contains the vertices of $G$ that are incident to the corresponding face. We modify this tree to consider the piercing edges, ensuring that the sizes of bags do not exceed $(c+7)/2$. (Recall that the width of a tree decomposition is the size of the largest bag minus 1.)
If $\mathcal{T}$ has an unpierced inner link $\lambda=\{u,v\}$, we split $\mathcal{T}$ along $\lambda$ into two triangulations and compute tree decompositions for them. Since each of them has a bag that contains $\lambda$, we can connect them and obtain a tree decomposition of $G$.
Hence, from now on, we assume that every inner link of $\mathcal{T}$ is pierced at least once. We root $T$ at an arbitrary leaf. The basic idea is as follows. Consider an edge $\{u,v\}$ that pierces inner links of $\mathcal{T}$. Then we lift both $u$ and $v$ to the bag $b$ that is the lowest common ancestor of the highest bags $b_{u}$ and $b_{v}$ that contain $u$ and $v$, respectively. That is, we place a copy of $u$ into each bag on the way from $b_{u}$ to $b$ and symmetrically for $v$. If we apply this strategy naively, then there may be a bag with $c+3$ vertices in the worst case. In the following, we describe how to improve this bound.
Consider a face $f$ of the triangulation $\mathcal{T}$. Let $b_{f}$ be the bag in $T$ that corresponds to $f$. Let $v_{1}$, $v_{2}$, and $v_{3}$ be the vertices of $f$ such that $f$ shares the link $\{v_{1},v_{2}\}$ with the face that corresponds to the parent of $f$ in $T$ (if any). An edge of $E(G)\setminus E(\mathcal{T})$ is lineal if one of its endpoints is in the bag $b_{f}$ or above and the other endpoint is in a bag below $b_{f}$ in the tree $T$. A long edge of $E(G)\setminus E(\mathcal{T})$ is bent if its endpoints are in different subtrees of $T$.
We perform the following two modifications of $T$ for each face $f$; see Fig. 5.
  • We subdivide each edge of $T$ between $b_{f}$ and its children (if any) by a new copy bag that is a copy of $b_{f}$. We call these bags primary copy bags.
  • If $f$ has two children in $T$, assume that piercing number of $\lambda=\{v_{1},v_{3}\}$ is less or equal to that of $\lambda^{\prime}=\{v_{2},v_{3}\}$ and that $\lambda$ is the common link of $f$ and its left child face. If there are $l>0$ bent edges $\{u_{1},w_{1}\},\dots,\{u_{l},w_{l}\}$ w.r.t. $f$, then we subdivide the edge of $T$ between $b_{f}$ and its left child by additional $l$ copies of $b_{f}$. These secondary copy bags are numbered $b_{f,1},\dots,b_{f,l}$ from top to bottom; see Fig. 5.
Figure 5: Steps 1 and 2 of the modification. Copy bags are oval; secondary copy bags are yellow.
Next, we lift vertices in bottom-up order, e.g., in preorder. More precisely, at each original bag $b_{f}$ (that is, not a copy bag) in our preorder traversal of $T$, we process each edge $e$ of $G$ that pierces $f$, according to its type as follows.
  • If $e$ is a short edge, that is, it is incident to a vertex of $f$, namely to $v_{1}$ or $v_{2}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ by one bag (as $v_{5}$ and $v_{7}$ in Fig. 6(a)) to the primary copy bag of $b_{f}$.
  • If $e$ is a long edge that pierces the link $\{v_{1},v_{2}\}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ all the way to $b_{f}$; see Fig. 6(b).
  • If $e=\{u_{i},w_{i}\}$ for some $i\in\{1,\dots,l\}$, then we lift both endpoints of $e$. First, we lift the highest occurrence of $u_{i}$ from the left subtree to $b_{f,i}$. Second, we lift the highest occurrence of $w_{i}$ in the right subtree of $b_{f}$ to $b_{f}$. Third, we copy $w_{i}$ into each bag from $b_{f}$ down to $b_{f,i}$; see Fig. 6(c). In this way, the two endpoints of $e$ meet in $b_{f,i}$.
Figure 6: Lifting up endpoints of the edges that pierce links of the face $f$.
The modifications described above make sure that the resulting tree $T^{\prime}$ is indeed a tree decomposition of $G$, i.e., for each vertex $v$ of $G$, the subtree induced by the bags that contain $v$ is connected, and for each edge $e$ of $G$, there is a bag in $T^{\prime}$ that contains both endpoints of $e$.
In the following, we show that each bag of $T^{\prime}$ contains at most $(c+7)/2$ vertices, i.e., we have added at most $(c+1)/2$ to the initial three vertices. For each face $f$, the number of vertices added to the original bag $b_{f}$ is the number of long edges w.r.t. $f$, which is at most $c/2$. It remains to bound the number of vertices that are added to each copy bag of $b_{f}$.
First, we show the bound for secondary copy bags. Recall that, for $i\in\{1,\dots,l\}$, we have added to $b_{f,i}$ from below $i$ endpoints of the bent long edges w.r.t. $f$ and from above $l+1-i$ endpoints of bent long edges; see Fig. 6(c). In addition, we have added a vertex for each lineal long edge. Therefore, the number of added vertices is at most one plus the number of long edges w.r.t. $f$ that pierce the link $\{v_{1},v_{3}\}$. Recall that the link $\{v_{1},v_{3}\}$ was chosen to have at most as many piercing edges as $\{v_{2},v_{3}\}$. Because each inner link of $\mathcal{T}$, including $\{v_{1},v_{2}\}$, is pierced at least once, $\{v_{1},v_{3}\}$ is pierced at most $(c-1)/2$ times. Thus, we have added at most $(c+1)/2$ vertices to $b_{f,i}$, and so $|b_{f,i}|\leq(c+7)/2$.
Now, we bound the number of vertices that are added to the primary (the bottommost) copy bag $b^{\prime}$ of $b_{f}$. Note that such a bag is a parent bag of some original bag $b_{g}$ such that the faces $f$ and $g$ are adjacent. There are two types of vertices that are added to $b^{\prime}$. The first type consists of the endpoints of the lineal long edges of $b_{g}$ (at most $c/2$). The second one is just one vertex $w$ of the face $g$ that is not shared with $f$ (for example, the endpoint $v_{5}$ of $g$ in Fig. 6(a)). We have added $w$ only if there are edges in $E(G)\setminus E(\mathcal{T})$ with lower endpoint $w$ (lower w.r.t. $T$). Each such edge contributes to the piercing number of $g$. Hence, the number of lineal long edges w.r.t. $g$ is at most $(c-1)/2$. In both cases ($w\in b^{\prime}$ and $w\not\in b^{\prime}$), we have added at most $(c+1)/2$ vertices to $b^{\prime}$, so $|b^{\prime}|\leq(c+7)/2$.
Lemmas 3.9 and 4.1 give us an upper bound $4$ for outer $2$-planar graphs, extending the known tight bound $k+2$ for $k=\{0,1\}$ [5, 6]. Note that $K_{5}$ is outer $2$-planar and $\operatorname{tw}(K_{5})=4$.
Theorem 4.3.
Every outer 2-planar graph has treewidth at most $4$, which is tight.
Theorem 4.4.
Every outer $k$-planar graph has treewidth at most $21.5k+$ 2.
Proof 4.5.
We already know better bounds for $k\leq 2$. For $k\geq 3$, by Lemmas 3.5, 4.1 and 3.7, we obtain a bound $2.51.5k+$ 2.5 for even $k$ and $21.5k+$ 2 for odd $k$. As treewidth is an integer, we obtain $21.5k+$ 2 for general $k$.
By Lemmas 3.11 and 4.1, we also obtain the following bound for outer min-$k$-planar graphs, which improves the previously known bound $113k+$ 11 by Wood and Telle [20]. Note that outer min-00-planar graphs are outerplanar graphs.
Theorem 4.6.
For $k\geq 1$, every outer min-$k$-planar graph has treewidth at most $13k+$ 1.
4.2 Separation Number
Theorems 4.4 and 4.6 immediately imply upper bounds on the separation number of outer $k$-planar and outer min-$k$-planar graphs. However, using our triangulations directly, we obtain even better bounds, namely $k+2$ for outer $k$-planar graphs and $12k+$ 1 for outer min-$k$-planar graphs. The first bound improves the bound of $32k+$ 3 by Chaplick et al. [10].
Lemma 4.7.
If $G$ is a graph with a convex drawing whose outer cycle admits a triangulation $\mathcal{T}$ with edge piercing number at most $c$, then $G$ has a balanced separator of size at most $c+2$.
Proof 4.8.
We construct a balanced separation of order at most $c+2$ using $\mathcal{T}$. In short, we select a link $\lambda$ of $\mathcal{T}$ that is a balanced separator for $\mathcal{T}$ and put its endpoints into the separator. Then, we add at most $c$ vertices to the separator according to the edges piercing $\lambda$.
First, we find a “centroid” triangle of $\mathcal{T}$ as follows. Let $T$ be the tree that is the weak dual of $\mathcal{T}$. It is well known that every tree contains a vertex such that, after removing it, the number of vertices in each subtree is at most half of the original tree. Let $f=\{u,v,w\}$ be the triangle corresponding to such a vertex of $T$. We partition $V\setminus f$ into three disjoint sets $V_{1},V_{2},V_{3}$; see Fig. 7(a). We may assume that $V_{1}$ is the largest among the three sets.
(a) “centroid” triangle $f$ and vertex sets $V_{1}$, $V_{2}$, $V_{3}$.
(b) edges piercing $\{u,v\}$ and separation $(A,B)$.
Figure 7: A triangulation with edge piercing number $c$ yields a balanced separator of size at most $c+2$.
Now, we construct the desired separation with the help of the sets $V_{1},V_{2},V_{3}$. As $\lambda=\{u,v\}$ is a link of $\mathcal{T}$, at most $c$ edges pierce $\lambda$, connecting vertices in $V_{1}$ with vertices in $V_{2}\cup\{w\}\cup V_{3}$. Let $S$ be the set of endpoints of the piercing edges on the latter side. Let $A=V_{1}\cup\{u,v\}\cup S$ and $B=V_{2}\cup\{u,v,w\}\cup V_{3}$; see Fig. 7(b). We claim that $(A,B)$ is a balanced separation of order at most $c+2$.
Clearly, the order of $(A,B)$ is $|A\cap B|=|S\cup\{u,v\}|$, which is at most $c+2$. Hence, it suffices to show that the sizes of $A\setminus B$ and $B\setminus A$ are at most $32n/$ 3. To this end, we first bound the size of $V_{1}$ as follows:
$\frac{n}{3}-1\leq|V_{1}|\leq\frac{n}{2}-1.$
The lower bound holds as $|V_{1}|+|V_{2}|+|V_{3}|=n-3$ and $|V_{1}|$ is the largest. The upper bound can be shown by the fact that a triangulated outerplanar graph has exactly $F+2$ vertices, where $F$ is the number of triangles. By the way of choosing $f$, the vertex set $V_{1}\cup\{u,v\}$ induces at most $n/2-1$ triangles. Therefore, $|V_{1}|=|V_{1}\cup\{u,v\}|-2\leq(n/2-1+2)-2=n/2-1$.
Now we can confirm that the sizes of $A\setminus B$ and $B\setminus A$ are at most $32n/$ 3 as follows:
$|A\setminus B|$ $=|V_{1}|\leq\frac{n}{2}-1\leq\frac{2}{3}n\text{\leavevmode\nobreak\ and}$
$|B\setminus A|$ $\leq|V_{2}\cup\{w\}\cup V_{3}|=|V\setminus(V_{1}\cup\{u,v\})|\leq n-\left(\frac{n}{3}-1\right)-2=\frac{2}{3}n-1.$
For every $k$, the classes of outer $k$-planar graphs and outer min-$k$-planar graphs are closed under taking subgraphs. Hence, by Lemmas 4.7, 3.5 and 3.11, we obtain the following bounds.
Theorem 4.9.
Every outer $k$-planar graph has separation number at most $k+2$, and for $k\geq 1$, every outer min-$k$-planar graph has separation number at most $12k+$ 1.
5 Lower Bounds
In this section, we complement the results in Section 4, by giving lower bounds for the separation number and treewidth of outer $k$-planar graphs. To show them, we use grid-like graphs called stacked prisms. The $m\times n$ stacked prism $Y_{m,n}$ is the (planar) graph obtained by connecting all pairs vertices in the topmost and bottommost row of the $m\times n$ grid that are in the same column; see Fig. 8(a).
(a) grid-like drawing
(b) clipping of a circular drawing
Figure 8: Two drawings of the stacked prism $Y_{m,n}$.
Theorem 5.1.
For every even number $k\geq 0$, there exists a graph $G$ such that $\operatorname{lcr}^{\circ}(G)=k$ and $\operatorname{sn}(G)=k+2$.
Proof 5.2.
We first observe that, if $m$ is even, $\operatorname{lcr}^{\circ}(Y_{m,n})$ is at most $22n-$ 2. If $m$ is even, we can place the $m$ rows of $Y_{m,n}$, each of length $n$, one after the other, alternating in their direction, around a circle; see Fig. 8(b). The grid has two types of edges: the row edges and the column edges, which connect vertices in the same row or column, respectively. As Fig. 8(b) shows, the row edges have no crossing. For each $i\in[n]$, there is a column edge $e_{i}$ that spans $22i-$ 2 other vertices along the perimeter of the circle. Each of these vertices is incident to exactly one column edge that crosses $e_{i}$. Hence, every column edge has at most $22n-$ 2 crossings, and $\operatorname{lcr}^{\circ}(Y_{m,n})\leq 2n-2$.
Let $n=k/2+1$, and let $m$ be a sufficiently large even number. Then we claim that $Y_{m,n}$ fulfills the conditions. As we discussed $\operatorname{lcr}^{\circ}(Y_{m,n})\leq k$ holds and therefore $\operatorname{sn}(Y_{m,n})\leq k+2$ follows from Theorem 4.9. Hence, it suffices to show $\operatorname{sn}(Y_{m,n})\geq k+2=2n$, which also implies $\operatorname{lcr}^{\circ}(Y_{m,n})\geq 2n-2=k$ by Theorem 4.9.
Suppose that $Y_{m,n}$ has a balanced separator $S$ of size less than $2n$. As there are $n$ columns, there is a column that contains at most one vertex of $S$. This column contains a path $P$ of length $m-1$ that does not intersect $S$. Now observe that at least $m-2n$ rows do not contain any vertex of $S$, and therefore, all vertices in these rows are connected to $P$. Hence, after removing $S$, the size of the connected component $C$ that contains $P$ is at least $n(m-2n)$. The ratio $|V(C)|/|V(Y_{m,n})|$ is $1-(2n/m)$, which is greater than $2/3$ if $m$ is sufficiently large.
Aidun et al. [1] showed that $\operatorname{tw}(Y_{m,n})=2n$ if $m>2n$. With the same stacked prism, we obtain the following lower bound.
Theorem 5.3.
For every even number $k\geq 0$, there exists a graph $G$ such that $\operatorname{lcr}^{\circ}(G)=k$ and $\operatorname{tw}(G)=k+2$.
6 Conclusion and Open Problems
We have introduced methods for triangulating drawings of outer $k$-planar graphs such that the triangulation edges cross few graph edges. These triangulations yield better bounds on treewidth and separation number of outer $k$-planar graphs. Our method is constructive; the corresponding treewidth decomposition and balanced separation can be computed efficiently. Via our triangulations, we improved the multiplicative constant in the upper bound on the treewidth of outer $k$-planar graphs from $3$ to $1.5$; we showed a lower bound of $1$. What is the correct multiplicative constant? Finding triangulations of outer $k$-planar graphs with lower triangle piercing number could be a step in this direction. It would also be interesting to find other graph classes that admit triangulations with low triangle piercing number.

### v2

@@
 keywords:
-treewidth, outer $k$-planar graphs, outer min-$k$-planar graphs, cop number, separation number
+treewidth, outerplanar graphs, outer $k$-planar graphs, outer min-$k$-planar graphs, cop number, separation number
 1 Introduction
@@
 Recently treewidth-based techniques have been successfully applied for convex drawings, that is, straight-line drawings where vertices lie on a circle.
-For example, Bannister and Eppstein [8] showed that if a graph admits a convex drawing with at most $k$ crossings in total, then its treewidth is bounded by a function of $k$.
+For example, Bannister and Eppstein [8] ([7]) showed that if a graph admits a convex drawing with at most $k$ crossings in total, then its treewidth is bounded by a function of $k$.
 Furthermore, they used this fact to give, for a fixed $k$, a linear-time algorithm (via extended monadic second-order logic and Courcelle’s theorem [13]) to decide whether a convex drawing with at most $k$ crossings exists for a given graph.
 Chaplick et al.
-[12] generalized this to bundled crossings (a crossing of two bundles of edges such that in each bundle the edges travel in parallel, which is counted as one crossing).
+[12] ([11]) generalized this to bundled crossings (a crossing of two bundles of edges such that in each bundle the edges travel in parallel, which is counted as one crossing).
 Another prominent class of graphs with convex drawings is the class of outer $k$-planar graphs, that is, graphs that admit a convex drawing where each edge is crossed at most $k$ times.
@@
 Auer et al.
-[5, 6] showed that maximal outer $1$-planar graphs are chordal graphs and, therefore, have treewidth at most $3$, which is tight.
+[5, 6] ([4]) showed that maximal outer $1$-planar graphs are chordal graphs and, therefore, have treewidth at most $3$, which is tight.
 Another parameter linked to the treewidth is the separation number of a graph.
@@
 Angelini et al.
-[2] showed that $2$-layer $k$-planar graphs have pathwidth at most $k+1$ and gave a lower bound of $(k+3)/2$ for every odd number $k$.
+[2] ([3]) showed that $2$-layer $k$-planar graphs have pathwidth at most $k+1$ and gave a lower bound of $(k+3)/2$ for every odd number $k$.
 Hence, their result can be considered as a smooth generalization of the caterpillar result.
@@
 The piercing number of a link is the number of edges of $G$ that pierce the link.
-In particular the piercing number of outer links is 0, and if a link coincides with an edge, then its piercing number is at most $k$ by the outer $k$-planarity of $\Gamma$.
+In particular, the piercing number of outer links is 0, and if a link coincides with an edge, then its piercing number is at most $k$ by the outer $k$-planarity of $\Gamma$.
 We define the edge piercing number of $\mathcal{T}$ to be the maximum piercing number of any link of $\mathcal{T}$.
@@
 Our splitting procedure starts by declaring the link $\{v_{1},v_{n}\}$ active and considers all other vertices to lie on the right side of the link.
-See Fig.
-1 for illustration.
+See Figure 1 for illustration.
 Clearly, the link $\{v_{1},v_{n}\}$ is not pierced.
@@
 We set $\varepsilon>0$ to a sufficiently small value.
-Then we place $v_{i}$ at $(0,-1)$, $v_{k}$ at $(0,1)$, all the vertices from the left side of the link within arc distance $\varepsilon$ of $(-1,0)$ and spread the vertices from the right side of the link evenly along the right semicircle, see Fig.
-2.
+Then we place $v_{i}$ at $(0,-1)$, $v_{k}$ at $(0,1)$, all the vertices from the left side of the link within arc distance $\varepsilon$ of $(-1,0)$ and spread the vertices from the right side of the link evenly along the right semicircle, see Figure 2.
 Figure 2: We can push crossings among the edges that pierce the active link $\{v_{i},v_{k}\}$ to the left.
@@
 As every pair of intertwined piercing edges has their left endpoints in arc distance at most $\varepsilon$ from $(-1,0)$ and their right endpoints in arc distance at least $\pi/n$ from each other, the crossing of two such segments is to the left of the active link.
-Lemma 3.1 gives us an equivalent drawing of $G$ in which piercing edges cross the active link in the same order from bottom to top as their right endpoints in the cyclic order.
+Lemma 3.1 gives us an equivalent drawing of $G$ in which piercing edges cross the active link in the same order from bottom to top as their right endpoints occur in the cyclic order.
 This allows us to draw the piercing edges without crossings on the right side of the active link.
@@
 We prove the lemma by providing an appropriate splitting procedure that recursively constructs the triangulation.
-Let $\Gamma$ be an outer $k$-planar drawing of $G$.
+Let $\Gamma$ be an outer $k$-planar drawing of the given maximal outer $k$-planar graph $G$.
 We prove by induction on the number of steps of the splitting procedure that every active link has piercing number at most $12k-$ 1.
@@
 In the first case, $\lambda$ is not pierced by any edge of $G$.
-Observe that $\{v_{i},v_{i+1}\}$ is an edge, so $u$ has at least one neighbor in the right side of $\lambda$ in $G$.
-We select the split vertex, $w$, to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$ in the circular order; see Fig.
-3(a).
+Observe that $\{v_{i},v_{i+1}\}$ is an edge, so $u$ has at least one neighbor on the right side of $\lambda$ in $G$.
+We select the split vertex, $w$, to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$ in the circular order; see Figure 3(a).
 In this case, the link $\{u,w\}$ coincides with an edge of $G$, and thus is pierced by at most $k$ edges.
@@
 Using Lemma 3.1, we modify the drawing $\Gamma$ so that the piercing edges cross $\lambda$ in the same bottom-to-top order as their right endpoints.
-Let $e_{1},e_{2},\dots,e_{\ell}$ denote the piercing edges and $w_{1},w_{2},\dots,w_{\ell}$ be their respective, not necessarily different, right endpoints in the order from bottom to top.
+Let $e_{1},e_{2},\dots,e_{\ell}$ denote the piercing edges, and let $w_{1},w_{2},\dots,w_{\ell}$ denote their respective, not necessarily different, right endpoints in the order from bottom to top.
 By induction, we have $\ell\leq 2k-1$.
@@
 (b) constructing the new triangle $\triangle$ bounded by $\lambda$, $c^{\prime}$, and $c^{\prime\prime}$; the orange edges that pierce both $c^{\prime}$ and $c^{\prime\prime}$ also cross the edge $e_{j}$.
-Figure 3: Constructing a triangulation $\mathcal{T}$ with edge piercing number $12k-$ 1.
-To bound the number of edges that pierce the link $\{u,w\}$, we draw the curve $c^{\prime}$ as in Fig.
-3(b).
+Figure 3: Constructing a triangulation $\mathcal{T}$ with edge piercing number at most $12k-$ 1.
+To bound the number of edges that pierce the link $\{u,w\}$, we draw the curve $c^{\prime}$ as in Figure 3(b).
 We start the curve $c^{\prime}$ at $u$ and follow $\lambda$ until before it crosses $e_{j}$, then follow $e_{j}$ until we arrive at $w$.
-The curve $c^{\prime}$ has crossings with at most $\lceil(\ell-1)/2\rceil$ edges that pierce $\lambda$, and at most $k$ edges that have a crossing with $e_{j}$.
+The curve $c^{\prime}$ crosses at most $\lceil(\ell-1)/2\rceil$ edges that pierce $\lambda$ and at most $k$ edges that pierce $e_{j}$.
 Thus, there are at most $\lceil(\ell-1)/2\rceil+k\leq\lceil(2k-2)/2\rceil+k=2k-1$ edges that pierce the link $\{u,w\}$.
-We can argue symmetrically for the link $\{w,v\}$; see curve $c^{\prime\prime}$ in Fig.
-3(b).
-As we have bounded the piercing number of every active link during the triangulation by $12k-$ 1, we get the desired bound of $12k-$ 1 for the edge piercing number of the constructed triangulation.
-Next, we present our refined triangulation method, where we consider an additional edge to get better bounds for the edge piercing number.
+We can argue symmetrically for the link $\{w,v\}$; see curve $c^{\prime\prime}$ in Figure 3(b).
+As we have bounded the piercing number of every active link during the triangulation procedure by $12k-$ 1, we get the desired bound of $12k-$ 1 for the edge piercing number of the constructed triangulation.
+Next, we present our refined triangulation method, where we consider an additional edge to get a better bound for the edge piercing number.
 Lemma 3.5.
@@
 Proof 3.6.
-The proof is similar to the proof of Lemma 3.3 and we use the same terminology.
+The proof is similar to the proof of Lemma 3.3, and we use the same terminology.
+Let again $\Gamma$ be an outer $k$-planar drawing of the given maximal outer $k$-planar graph $G$.
 We present a splitting procedure and show that every active link is pierced by at most $k$ edges.
-For the base of induction, the link $\{v_{1},v_{n}\}$ has no piercing edges.
+For the base of the induction, observe that the link $\{v_{1},v_{n}\}$ has no piercing edges.
 Now, let $\lambda=\{v_{i},v_{k}\}$ be the active link and set $u=v_{i}$ and $v=v_{k}$.
-If $\lambda$ has no piercing edges, we proceed as in the proof of Lemma 3.3 and pick the split vertex to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$.
+If $\lambda$ has no piercing edges in $\Gamma$, we proceed as in the proof of Lemma 3.3 and pick the split vertex to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$.
 Otherwise, $\lambda$ is pierced by edges of $G$.
 Using Lemma 3.1, we assume that the piercing edges cross the active link in the same bottom-to-top order as their right endpoints.
-Now, let $e_{1},e_{2},\ldots,e_{\ell}$ be the piercing edges.
+Now, let $e_{1},e_{2},\ldots,e_{\ell}$ be the piercing edges, and let $w_{1},w_{2},\dots,w_{\ell}$ be their respective, not necessarily different, right endpoints in the order from bottom to top.
 We pick the middle edge $e_{j}$ with $j=\lceil(\ell+1)/2\rceil$.
-If $e_{j}$ has no more crossings on the right of the active link, then we pick the right endpoint of $e_{j}$ to be the split vertex.
-Both curves that start in the right endpoint of $e_{j}$, follow $e_{j}$ to $e$, and then follow $\lambda$ to either $u$ or $v$ have at most $\lceil(\ell-1)/2\rceil\leq k$ crossings each, so we can proceed.
-Now, let $x$ be the first crossing on the edge $e_{j}$ that occurs on the right side of $\lambda$, and let $\hat{e}$ denote the edge that crosses $e_{j}$ at $x$; see Fig.
-4(a).
+If $e_{j}$ has no crossings on the right side of the active link, then we pick $w_{j}$ to be the split vertex.
+The two curves that start in $w_{j}$, follow $e_{j}$ to $e$, and then follow $\lambda$ to either $u$ or $v$ have at most $\lceil(\ell-1)/2\rceil\leq k$ crossings each, so we can proceed.
+Now, let $x$ denote the first crossing on the edge $e_{j}$ that occurs on the right side of $\lambda$, and let $\hat{e}$ denote the edge that crosses $e_{j}$ at $x$; see Figure 4(a).
+(a) $\lambda$ and $\hat{e}$ do not share an endpoint; the new triangle $\triangle$ is bounded by $\lambda$, $c^{\prime}$, and $c^{\prime\prime}$.
+(b) $\lambda$ and $\hat{e}$ share an endpoint; the new triangle $\triangle$ is bounded by $\lambda$, $\hat{e}$, and $c^{\prime\prime}$.
+Figure 4: Constructing a triangulation $\mathcal{T}$ with edge piercing number $k$.
 We first consider the case where $\hat{e}$ and $\lambda$ are disjoint.
@@
 The case $u<w<w_{j}$ is symmetric.
-(a) $\lambda$ and $\hat{e}$ do not share an endpoint; the new triangle $\triangle$ is bounded by $\lambda$, $c^{\prime}$, and $c^{\prime\prime}$.
-(b) $\lambda$ and $\hat{e}$ share an endpoint; the new triangle $\triangle$ is bounded by $\lambda$, $\hat{e}$, and $c^{\prime\prime}$.
-Figure 4: Constructing a triangulation $\mathcal{T}$ with edge piercing number $k$.
 We bound the piercing number of the link $\{v,w\}$.
-We start the curve $c^{\prime\prime}$ at $v$ and follow the link $\lambda$ until just before it crosses $e_{j}$, then follow $e_{j}$ until just before it crosses $\hat{e}$, then follow $\hat{e}$ until we arrive at $w$; see Fig.
-4(a).
+To this end, we start tracing a curve $c^{\prime\prime}$ at $v$ and follow the link $\lambda$ until just before it crosses $e_{j}$, then we follow $e_{j}$ until just before it crosses $\hat{e}$, and then we follow $\hat{e}$ until we arrive at $w$; see Figure 4(a).
 In the first part, the curve crosses at most $\lceil(\ell-1)/2\rceil\leq\lceil(k-1)/2\rceil$ edges.
@@
 Thus, by Section 2, the piercing number of the link $\{v,w\}$ is at most $\lceil(k-1)/2\rceil+\lfloor(k-1)/2\rfloor\leq k-1$.
-To bound the piercing number of the link $\{u,w\}$, we start the curve $c^{\prime}$ at $u$ and follow the link $\lambda$ until just after it crosses $e_{j}$ and then follow the curve $c^{\prime\prime}$ to $w$, see Fig.
-4(a).
-This time, the number of crossings on the curve $c^{\prime}$ is at most $k$.
+To bound the piercing number of the link $\{u,w\}$, we start tracing a curve $c^{\prime}$ at $u$, follow the link $\lambda$ until just after it crosses $e_{j}$ and then follow the curve $c^{\prime\prime}$ to $w$, see Figure 4(a).
+The curve $c^{\prime}$ crosses the same set of edges as the curve $c^{\prime\prime}$ plus the edge $e_{j}$, so $c^{\prime\prime}$ has at most $k$ crossings.
 Thus, the piercing number of the link $\{u,w\}$ is at most $k$.
-Now we consider the last case, where $\hat{e}$ and $\lambda$ have a common vertex.
-In this case, we can choose the other vertex of $\hat{e}$ as $w$; see Fig.
-4(b).
-Without loss of generality, we assume that $u$ is the common vertex.
-The edge $\{u,w\}$ has at most $k$ crossings by $k$-planarity.
-To bound the piercing number of the link $\{v,w\}$, observe that the edge $\{u,w\}$ has exactly $j$ crossings on the curve from $v$ to $x$ (including $x$).
-Therefore, as we did in the above, the edge $\{v,w\}$ is crossed by at most $(\ell-j)+(k-j)=k-(2j-\ell)$ edges.
-Note that $2j-\ell$ is $1$ if $\ell$ is $odd$ and 00 otherwise.
+Now we consider the last case, where $\hat{e}$ and $\lambda$ have a common endpoint.
+In this case, we choose the other endpoint of $\hat{e}$ as $w$; see Figure 4(b).
+We first assume that $u$ is the common endpoint.
+The edge $\{u,w\}$ is an edge of $G$, so it has at most $k$ crossings.
+To bound the piercing number of the link $\{v,w\}$, we argue with the curve $c^{\prime\prime}$ defined in the previous case.
+Observe that the edge $\{u,w\}$ has exactly $j$ crossings on the part from $u$ to $x$ (including $x$).
+Therefore, we can argue as above and conclude that the edge $\{v,w\}$ is crossed by at most $(\ell-j)+(k-j)=k-(2j-\ell)$ edges.
+Note that $2j-\ell=2\lceil(\ell+1)/2\rceil-\ell$ is $1$ if $\ell$ is odd and $2$ otherwise.
+Similarly, if $v$ is the common endpoint, we obtain the bound $(j-1)+(k-(\ell-j+1))=k+(2j-\ell)-2$ for the piercing number of the link $\{u,w\}$.
 The piercing number of a triangle of $\mathcal{T}$ is the sum of the piercing numbers of the links that form the sides of the triangle.
@@
 Hence, the triangle piercing number is at most $4$.
-Lastly, we combine the two strategies from Lemmas 3.3 and 3.5 to obtain a triangulation method for outer min-$k$-planar graphs.
+Lastly, we combine the two triangulation strategies from the proofs of Lemmas 3.3 and 3.5 to obtain a triangulation method for outer min-$k$-planar graphs.
 Lemma 3.11.
@@
 We proceed as in the proof of Lemma 3.3.
-We want to give a splitting procedure for which every active link has piercing number at most $12k-$ 1.
-Consult Fig.
-3 for a reminder of the notation and the strategy.
-As $\Gamma$ is now an outer min-$k$-planar drawing, edges may have more than $k$ crossings.
-We call such edges heavy.
-Other edges, i.e., edges with no more than $k$ crossings, are light.
+We describe a splitting procedure such that every active link has piercing number at most $12k-$ 1.
+Consult Figure 3 for a reminder of the notation and the strategy.
+As the given drawing is outer min-$k$-planar, edges may now have more than $k$ crossings.
+We call such edges heavy and the other edges light.
 By definition, heavy edges cross only light edges.
-The first thing to notice is that as we allow for heavy edges now, the previous proof breaks in both cases when the curves $c^{\prime}$, $c^{\prime\prime}$ are routed along heavy edges.
-In particular, in the first case, $\lambda$ has no piercing edges, and we select the split vertex $w$ to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$.
-If edge $\{u,w\}$ is heavy, then we cannot bound the number of crossings along the curve $c^{\prime\prime}$.
-In the second case, we select the split vertex to be the right endpoint $w_{j}$ of the middle piercing edge $e_{j}$.
-If edge $e_{j}$ is heavy, then we cannot bound the numbers of crossings along $c^{\prime}$ and $c^{\prime\prime}$.
+We first explain why the previous proof breaks in both cases if the curves $c^{\prime}$ and $c^{\prime\prime}$ are routed along heavy edges.
+Recall that in the first case when $\lambda$ has no piercing edges, we selected the split vertex $w$ to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$.
+If, however, the edge $\{u,w\}$ is heavy, then we cannot bound the number of crossings along the curve $c^{\prime\prime}$.
+In the second case, we selected the split vertex to be the right endpoint $w_{j}$ of the middle piercing edge $e_{j}$.
+If the edge $e_{j}$ is heavy, then we cannot bound the numbers of crossings along $c^{\prime}$ and $c^{\prime\prime}$.
 To resolve the issue with the second case, we apply the strategy from the proof of Lemma 3.5.
-Observe that if $e_{j}$ is heavy, then for the first crossing $x$ with an edge $\hat{e}$ on the right side of $\lambda$, $\hat{e}$ is a light edge.
-Similarly, as in the proof of Lemma 3.5, see Fig.
-4, we select the split vertex $w$ to be the endpoint of $\hat{e}$ that is different than $u$ and $v$ and in case both endpoints are, the one on the side of $x$ with fewer crossings.
-The piercing numbers of new links are at most $\lceil(2k-1)/2\rceil+k-1\leq 2k-1$.
+Observe that if $e_{j}$ is heavy, then for the first crossing $x$ with an edge $\hat{e}$ on the right side of $\lambda$, the edge $\hat{e}$ is a light.
+Similarly, as in the proof of Lemma 3.5 (see Figure 4), we select the split vertex $w$ to be the endpoint of $\hat{e}$ that is different from $u$ and $v$ and in case both endpoints are, the one on the side of $x$ with fewer crossings.
+The piercing numbers of the two new links are at most $\lceil(2k-1)/2\rceil+k-1\leq 2k-1$.
 To resolve the issue with the first case, we use a similar strategy.
-When edge $\{u,w\}$ is heavy, let $x$ to be the first crossing on $\{u,w\}$ from $u$ with an edge $\hat{e}$.
-Notice that $\hat{e}$ is light.
+If the edge $\{u,w\}$ is heavy, let $x$ be the crossing on $\{u,w\}$ that is closest to $u$ and let $\hat{e}$ be the crossing edge.
+Note that $\hat{e}$ is light.
 We select the split vertex $w$ to be any right endpoint of $\hat{e}$.
-The curves that start at $w$, go along $\hat{e}$ to $x$, then along $\{u,w\}$ to $u$ and optionally to $v$ along $\lambda$ have at most $k$ crossings, only with edges that also cross $\hat{e}$.
+Now consider the two curves that start at $w$, go along $\hat{e}$ to $x$, then along $\{u,w\}$ to $u$, and optionally to $v$ along $\lambda$.
+Both curves have at most $k$ crossings, only with edges that also cross $\hat{e}$.
 Remark 3.13.
-Each step of the splitting procedures used in Lemmas 3.33.11 can be implemented to run in $O(k)$ time, and the corresponding triangulations can be constructed in $O(nk)$ time.
+Given the intersection graph of the edges of $G$, the triangulations in Lemmas 3.33.11 can be constructed in $O(nk)$ time.
+Each step of the splitting procedures presented in Lemmas 3.33.11 can be implemented to run in $O(k)$ time, as the edges that pierce the new active links also pierce at least one of the edges $\lambda$, $e_{j}$, and $\hat{e}$.
 4 Applications
@@
 A long edge of $E(G)\setminus E(\mathcal{T})$ is bent if its endpoints are in different subtrees of $T$.
-We perform the following two modifications of $T$ for each face $f$; see Fig. 5.
+We perform the following two modifications of $T$ for each face $f$; see Figure 5.
   • We subdivide each edge of $T$ between $b_{f}$ and its children (if any) by a new copy bag that is a copy of $b_{f}$. We call these bags primary copy bags.
@@
 $f$, then we subdivide the edge of $T$ between $b_{f}$ and its left child by additional $l$ copies of $b_{f}$.
-These secondary copy bags are numbered $b_{f,1},\dots,b_{f,l}$ from top to bottom; see Fig.
-5.
+These secondary copy bags are numbered $b_{f,1},\dots,b_{f,l}$ from top to bottom; see Figure 5.
 Figure 5: Steps 1 and 2 of the modification. Copy bags are oval; secondary copy bags are yellow.
@@
 More precisely, at each original bag $b_{f}$ (that is, not a copy bag) in our preorder traversal of $T$, we process each edge $e$ of $G$ that pierces $f$, according to its type as follows.
-  • If $e$ is a short edge, that is, it is incident to a vertex of $f$, namely to $v_{1}$ or $v_{2}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ by one bag (as $v_{5}$ and $v_{7}$ in Fig.
-6(a)) to the primary copy bag of $b_{f}$.
-  • If $e$ is a long edge that pierces the link $\{v_{1},v_{2}\}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ all the way to $b_{f}$; see Fig. 6(b).
+  • If $e$ is a short edge, that is, it is incident to a vertex of $f$, namely to $v_{1}$ or $v_{2}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ by one bag (as $v_{5}$ and $v_{7}$ in Figure 6(a)) to the primary copy bag of $b_{f}$.
+  • If $e$ is a long edge that pierces the link $\{v_{1},v_{2}\}$, then we lift the highest occurrence of the other endpoint $u$ of $e$ all the way to $b_{f}$; see Figure 6(b).
   • If $e=\{u_{i},w_{i}\}$ for some $i\in\{1,\dots,l\}$, then we lift both endpoints of $e$.
@@
 Second, we lift the highest occurrence of $w_{i}$ in the right subtree of $b_{f}$ to $b_{f}$.
-Third, we copy $w_{i}$ into each bag from $b_{f}$ down to $b_{f,i}$; see Fig.
-6(c).
+Third, we copy $w_{i}$ into each bag from $b_{f}$ down to $b_{f,i}$; see Figure 6(c).
 In this way, the two endpoints of $e$ meet in $b_{f,i}$.
@@
 Recall that, for $i\in\{1,\dots,l\}$, we have added to $b_{f,i}$ from below $i$ endpoints of the bent long edges w.r.t.
-$f$ and from above $l+1-i$ endpoints of bent long edges; see Fig.
-6(c).
+$f$ and from above $l+1-i$ endpoints of bent long edges; see Figure 6(c).
 In addition, we have added a vertex for each lineal long edge.
@@
 The first type consists of the endpoints of the lineal long edges of $b_{g}$ (at most $c/2$).
-The second one is just one vertex $w$ of the face $g$ that is not shared with $f$ (for example, the endpoint $v_{5}$ of $g$ in Fig.
-6(a)).
+The second one is just one vertex $w$ of the face $g$ that is not shared with $f$ (for example, the endpoint $v_{5}$ of $g$ in Figure 6(a)).
 We have added $w$ only if there are edges in $E(G)\setminus E(\mathcal{T})$ with lower endpoint $w$ (lower w.r.t.
@@
 Let $f=\{u,v,w\}$ be the triangle corresponding to such a vertex of $T$.
-We partition $V\setminus f$ into three disjoint sets $V_{1},V_{2},V_{3}$; see Fig.
-7(a).
+We partition $V\setminus f$ into three disjoint sets $V_{1},V_{2},V_{3}$; see Figure 7(a).
 We may assume that $V_{1}$ is the largest among the three sets.
@@
 Let $S$ be the set of endpoints of the piercing edges on the latter side.
-Let $A=V_{1}\cup\{u,v\}\cup S$ and $B=V_{2}\cup\{u,v,w\}\cup V_{3}$; see Fig.
-7(b).
+Let $A=V_{1}\cup\{u,v\}\cup S$ and $B=V_{2}\cup\{u,v,w\}\cup V_{3}$; see Figure 7(b).
 We claim that $(A,B)$ is a balanced separation of order at most $c+2$.
@@
 To show them, we use grid-like graphs called stacked prisms.
-The $m\times n$ stacked prism $Y_{m,n}$ is the (planar) graph obtained by connecting all pairs vertices in the topmost and bottommost row of the $m\times n$ grid that are in the same column; see Fig.
-8(a).
+The $m\times n$ stacked prism $Y_{m,n}$ is the (planar) graph obtained by connecting all pairs vertices in the topmost and bottommost row of the $m\times n$ grid that are in the same column; see Figure 8(a).
 (a) grid-like drawing
@@
 We first observe that, if $m$ is even, $\operatorname{lcr}^{\circ}(Y_{m,n})$ is at most $22n-$ 2.
-If $m$ is even, we can place the $m$ rows of $Y_{m,n}$, each of length $n$, one after the other, alternating in their direction, around a circle; see Fig.
-8(b).
+If $m$ is even, we can place the $m$ rows of $Y_{m,n}$, each of length $n$, one after the other, alternating in their direction, around a circle; see Figure 8(b).
 The grid has two types of edges: the row edges and the column edges, which connect vertices in the same row or column, respectively.
-As Fig.
-8(b) shows, the row edges have no crossing.
+As Figure 8(b) shows, the row edges have no crossing.
 For each $i\in[n]$, there is a column edge $e_{i}$ that spans $22i-$ 2 other vertices along the perimeter of the circle.

### v3

@@
 We set $\varepsilon>0$ to a sufficiently small value.
-Then we place $v_{i}$ at $(0,-1)$, $v_{k}$ at $(0,1)$, all the vertices from the left side of the link within arc distance $\varepsilon$ of $(-1,0)$ and spread the vertices from the right side of the link evenly along the right semicircle, see Figure 2.
+Then we place $v_{i}$ at $(0,-1)$, $v_{k}$ at $(0,1)$, all the vertices on the left side of the link within arc distance $\varepsilon$ of $(-1,0)$ and spread the vertices from the right side of the link evenly along the right semicircle, see Figure 2.
 Figure 2: We can push crossings among the edges that pierce the active link $\{v_{i},v_{k}\}$ to the left.
@@
 (a) $\lambda$ and $\hat{e}$ do not share an endpoint; the new triangle $\triangle$ is bounded by $\lambda$, $c^{\prime}$, and $c^{\prime\prime}$.
-(b) $\lambda$ and $\hat{e}$ share an endpoint; the new triangle $\triangle$ is bounded by $\lambda$, $\hat{e}$, and $c^{\prime\prime}$.
+(b) $\lambda$ and $\hat{e}$ share the endpoint $u$; the new triangle $\triangle$ is bounded by $\lambda$, $\hat{e}$, and $c^{\prime\prime}$.
 Figure 4: Constructing a triangulation $\mathcal{T}$ with edge piercing number $k$.
@@
 Thus, the piercing number of the link $\{u,w\}$ is at most $k$.
-Now we consider the last case, where $\hat{e}$ and $\lambda$ have a common endpoint.
-In this case, we choose the other endpoint of $\hat{e}$ as $w$; see Figure 4(b).
+Now we consider the last case, where $\hat{e}$ and $\lambda$ have a common endpoint. In this case, we choose the other endpoint of $\hat{e}$ as $w$ (see Figure 4(b)) and analyse the two subcases.
 We first assume that $u$ is the common endpoint.
-The edge $\{u,w\}$ is an edge of $G$, so it has at most $k$ crossings.
-To bound the piercing number of the link $\{v,w\}$, we argue with the curve $c^{\prime\prime}$ defined in the previous case.
+Then the link $\{u,w\}=\hat{e}$ is an edge of $G$, so it has at most $k$ crossings.
+To bound the piercing number of the link $\{v,w\}$, we argue with the curve $c^{\prime\prime}$ that goes from $v$ to $w$ as defined in the previous case.
 Observe that the edge $\{u,w\}$ has exactly $j$ crossings on the part from $u$ to $x$ (including $x$).
-Therefore, we can argue as above and conclude that the edge $\{v,w\}$ is crossed by at most $(\ell-j)+(k-j)=k-(2j-\ell)$ edges.
-Note that $2j-\ell=2\lceil(\ell+1)/2\rceil-\ell$ is $1$ if $\ell$ is odd and $2$ otherwise.
-Similarly, if $v$ is the common endpoint, we obtain the bound $(j-1)+(k-(\ell-j+1))=k+(2j-\ell)-2$ for the piercing number of the link $\{u,w\}$.
+Therefore, we can argue as before and conclude that the piercing number of the edge $\{v,w\}$ is at most $(\ell-j)+(k-j)=k-(2j-\ell)$.
+This is bounded from above by $k-1$ since $2j-\ell=2\lceil(\ell+1)/2\rceil-\ell$ is $1$ if $\ell$ is odd and $2$ otherwise.
+Similarly, if $v$ is the common endpoint, then the link $\{v,w\}=\hat{e}$ is an edge of $G$, so it has at most $k$ edges, and the piercing number of the link $\{u,w\}$ is at most $(j-1)+(k-(\ell-j+1))=k+(2j-\ell)-2\leq k$.
 The piercing number of a triangle of $\mathcal{T}$ is the sum of the piercing numbers of the links that form the sides of the triangle.
@@
 If $\ell$ is even, it implies $\ell\neq k$ and therefore the link $\lambda$ is pierced by $\ell\leq k-1$ edges.
-Slightly improving the analysis in the proof of Lemma 3.5 for $k=2$ yields the following.
+With a slight modification of the strategy described in the proof of Lemma 3.5 and a careful analysis, we also obtain a better bound of the triangle piercing number for the case $k=2$.
 Lemma 3.9.
 Every outer $2$-planar drawing of a maximal outer $2$-planar graph admits a triangulation of the outer cycle with edge piercing number at most $2$ and triangle piercing number at most $4$.
+Figure 5: $\lambda$ and $\hat{e}$ share the endpoint $v$; the new triangle $\triangle$ is bounded by $\lambda$, $c^{\prime}$, and $c^{\prime\prime}$.
 Proof 3.10.
-We bound the triangle piercing number more carefully.
-If $\lambda$ has no piercing edges or $e_{j}$ has no crossings on the right side, then the claim is clearly true.
-If $e_{j}$ has a crossing, the curves $c^{\prime}$ and $c^{\prime\prime}$ meet each other near $e_{j}$ with exactly $l$ crossings in total and then proceed to $w$ without crossings.
-Hence, the triangle piercing number is at most $4$.
+We roughly follow the strategy of the proof of Lemma 3.5.
+Recall that $\lambda$ is the active link and that $\ell$ is the number of edges that pierce $\lambda$.
+If $\ell>0$, then $e_{j}$ with $j=\lceil(\ell+1)/2\rceil$ is the “middle” edge among those that pierce $\lambda$.
+If $e_{j}$ has crossings on the right side of $\lambda$, then, among these, $x$ is the crossing point of $e_{j}$ closest to $\lambda$.
+The edge $\hat{e}$ is the edge that crosses $e_{j}$ in $x$.
+As in the proof of Lemma 3.5, we distinguish the following cases.
+  1. 1.
+    $\lambda$ has no piercing edges.
+  2. 2.
+    $e_{j}$ has no crossings to the right of $\lambda$.
+  3. 3.
+    $\lambda$ and $\hat{e}$ are disjoint (Figure 4(a)).
+  4. 4.
+    $\lambda$ and $\hat{e}$ meet in $u$ (Figure 4(b)).
+  5. 5.
+    $\lambda$ and $\hat{e}$ meet in $v$.
+In cases 14, we choose the split vertex $w$ as in the proof of Lemma 3.5, which means that the new active links are pierced at most $k=2$ times.
+In case 5, we now choose $w$ to be $w_{j}$ (see Figure 5).
+Let $c^{\prime}$ be a curve that starts in $u$, follows $\lambda$ until just before $e_{j}$ and then follows $e_{j}$ to $w$.
+Let $c^{\prime\prime}$ be a curve that starts in $v$, follows $\hat{e}$ to $x$, and then follows $e_{j}$ to $w$.
+In order to show that each of the two curves is crossed by at most two edges, we need to do some preparations.
+We say that an edge lies to the left of the active link if both of its endpoints lie on the left side or one endpoint lies on the left and the other is an endpoint of the active link.
+We further say that an edge $e$ piercing a link $\lambda$ is anchored with respect to $\lambda$ if there exists another edge $f$ that crosses $e$ and either $f=\lambda$ or $f$ lies to the left of $\lambda$.
+Observe that an anchored edge can have at most one crossing to the right of $\lambda$.
+We claim that, during the splitting procedure described above, every edge piercing an active link is anchored (with respect to the active link).
+The claim is true in the beginning as $\lambda=\{v_{1},v_{n}\}$ is not pierced.
+In case 1, every edge that pierces $\{u,w\}$ or $\{v,w\}$ also crosses the edge $\{u,w\}$, so it is anchored.
+In case 2, every edge that pierces $\{u,w\}$ or $\{v,w\}$ also pierces $\lambda=\{u,v\}$ and is anchored by induction.
+In case 3, observe that no edge crosses $e_{j}$ between $\lambda$ and $x$ (as $x$ is the first crossing on $e_{j}$ to the right of $\lambda$), and no edge crosses $\hat{e}$ between $x$ and $w$ (by the choice of $w$ and since $\lfloor(k-1)/2\rfloor=0$ for $k=2$).
+Therefore, every edge that pierces $\{u,w\}$ or $\{v,w\}$ also pierces $\lambda$ and is anchored by induction.
+In case 4, $\{u,w\}$ is an edge, so every edge piercing it is anchored.
+Every edge that pierces $\{v,w\}$ either crosses $\{u,w\}$ or pierces $\lambda$ and is anchored by induction.
+Similarly, in case 5, every edge that pierces $\{u,w\}$ or $\{v,w\}$ either pierces $\lambda$ or crosses $e_{j}$ (which lies to the left of both new active links).
+We now return to case 5 and bound the piercing numbers of the new active links $\{u,w\}$ and $\{v,w\}$.
+To do this, we count the edges that cross the curves $c^{\prime}$ and $c^{\prime\prime}$, respectively.
+As $e_{j}$ is anchored with respect to $\lambda$, we get that $x$ is the only crossing on $e_{j}$ to the right of $\lambda$.
+Now, it is easy to see that $c^{\prime}$ is pierced only by $\hat{e}$ and possibly by $e_{1}$ in case $j=2$.
+For $c^{\prime\prime}$, we can show that there are no crossings along the curve.
+First, observe that any edge that crosses $\hat{e}$ between $v$ and $x$ has to pierce $\lambda$, which is not possible.
+Second, no edge crosses $e_{j}$ between $x$ and $w$, as $e_{j}$ is anchored and $x$ is the only crossing on $e_{j}$ to the right of $\lambda$.
+Combining these two observations, we see that there is no place for $c^{\prime\prime}$ to cross with an edge.
+It remains to show that the triangle piercing number of the new triangle $\triangle$ formed by $u$, $v$, and $w$ is at most $4$.
+In case 1, $\lambda$ has no piercing edges.
+In cases 2, 4, and 5, $\{v,w\}$ has no piercing edges.
+In case 3, we distinguish two subcases depending on the relative position of $w$.
+When $w_{j}<w$, then $\{v,w\}$ has no piercing edges.
+When $w<w_{j}$, then each of $\{u,w\}$ and $\{v,w\}$ is pierced at most once.
 Lastly, we combine the two triangulation strategies from the proofs of Lemmas 3.3 and 3.5 to obtain a triangulation method for outer min-$k$-planar graphs.